# Shot noise in Aharonov-Bohm interferometers: Comparison of helical and conventional setups

### v1

@@ Abstract
We study the shot noise of current through the edge states of a two-dimensional topological insulator placed in magnetic field and compare the obtained results with the shot noise in conventional single-channel spinless Aharonov-Bohm interferometer. We find general formulas for the Fano factors of these setups, assuming that temperature exceeds level spacing in the system. We demonstrate that both in helical and in conventional case the interference effects dramatically change the Fano factor and its magnetic field dependence. For weak tunneling coupling with leads, the Fano factors of both setups exhibit a periodic series of sharp Aharonov-Bohm peaks with variation of magnetic flux piercing the system. Our key finding is that the Fano factor in the helical interferometer provides information about the presence of backscattering defects violating topological protection. In particular, the amplitude of Aharonov-Bohm peaks in the helical setup is proportional to the strength of the defect in contrast to conventional setup, where peaks have finite amplitude even in the ballistic case.
Helical Edge States, Aharonov-Bohm interferometry, Shot noise
@@ I Introduction
Quantum interferometry is a rapidly growing area of electronics with a huge potential for practical use, in particular, in the booming fields of quantum computing and quantum cryptography. One beautiful manifestation of quantum interference is the Aharonov-Bohm effect Aharonov and Bohm (1959, 1963)—the sensitivity of the phase of electron wave function to magnetic flux. Quantum interferometers based on this effect are used for high-precision measurements of magnetic fields, for example, by using SQUID circuits.
@@
The simplest setup to study interference and particularly the Aharonov-Bohm effect is a ring-shaped interferometer with two single-channel arms placed in the perpendicular magnetic field. Due to the interference of electron trajectories winding around such a setup, the observables, like transmission coefficient and the noise intensity, show Aahronov-Bohm oscillations with a certain perios $\Delta\phi$ as a function of dimensionless magnetic flux $\phi=\Phi/\Phi_{0}$ threading the device, where $\Phi_{0}=hc/e$ is the flux quantum. The key statement of our paper is that noise in such a simple system is still not well understood and depends on specific realizations of the single-channel arms.
@@
Below we will investigate and compare in details two types of setups: (i) a spinless single-channel interferometer based on conventional materials (CI), i.e. a single-channel spinless quantum wire, folded in a ring-like shape and having two tunnel or metallic contacts attached (see Fig. 1a); (ii) an interferometer based on helical edge states (HI). The latter interferometer can be implemented in two-dimensional topological insulators  Bernevig and Hughes (2013), which are insulators in the bulk, but have helical one-dimensional (1D) edge states conducting the current without dissipation Bernevig and Hughes (2013); Hasan and Kane (2010); Qi and Zhang (2011). These edge states are formed by electrons with opposite spins travelling in opposite directions, so that backscattering by non-magnetic impurities is prohibited. If one attaches two contacts to the edge and apply the gate voltage to shift the Fermi level into the band gap, then the conductance of such a device is fully determined by the edge states (see Fig. 1b). Since the sample boundary can be bypassed in two directions, possibly with several windings, such a system represents an interferometer that encodes information about properties of the helical edge states (HES).
Both setups, CI and HI, have two electron waves propagating clockwise and counterclockwise in 1D channel. However, the interference phenomena in these systems manifest themselves in a rather different way. Most interestingly, in case of ballistic HI, the waves propagating in opposite directions have also opposite spin projections and cannot interfere provided that contacts to the HES are non-magnetic. Interference effects in HI appear in the presence of backscattering. This means that the observation of interference phenomena in the HI can give information about mechanisms of possible violation of the topological protection. By contrast, a spinless CI based on conventional materials (Fig. 1a) shows interference effects even in the ballistic case, because non-magnetic contacts in such setup lead to backscattering. Hence, ballistic CI is analogous to HI with a magnetic defect (or with one or two magnetic contacts), see Fig. 1b. We will also demonstrate that $\Delta\phi$ is different in these cases: $\Delta\phi=1$ for CI and $\Delta\phi=1/2$ for HI with a single static magnetic defect.
@@
Figure 1: Aharonov-Bohm interferometers based on conventional spinless single-channel quantum wire (a) and on helical edge channels of 2D topological insulator (b). The leads (shown by grey color) are modeled by conventional single-channel [spinless (a) and spinfull (b)] wires. The black dot in panel (b) represents a magnetic impurity. Grey regions represent contacts described by $3\times 3$ and $4\times 4$ $S-$matrix for CI and HI, respectively [see Eqs. (2), (3) and Figs. 2, 4].
@@
For both setups we will investigate the manifestations of quantum interference in the shot noise of the current. Such noise is a consequence of the discreteness of the electron charge. Importantly, its measurement provides information not available from conductance measurements: namely, the charge and statistics of current carriers, and the internal energy scales of the system de Jong and Beenakker (1997); Blanter and Büttiker (2000). Our primary goal is to calculate the Fano factor, $\mathcal{F}$, which is the ratio of the shot noise and the so-called Schottky noise value. This ratio is often used to characterize different transport regimes. For example, the values $\mathcal{F}=1/3~{}$ and $\mathcal{F}=1/2~{}$ for a diffusive conductor and tunneling transmission through a quantum level, respectively, are universal and independent of the system details de Jong and Beenakker (1997); Blanter and Büttiker (2000).
@@
Since the conductance and shot noise of interferometers of various types were studied in a great number of works, we will start with a short description of the state of the art. Let us first recall the key results known for the conductance of the Aharonov-Bohm interferometer. At weak tunnel coupling and at low temperatures, the conductance $G(\phi)$ exhibits sharp resonances at certain values of magnetic flux, arising when quantum levels of the system are crossed by the Fermi level Büttiker et al. (1984, 1985). The case of relatively high temperatures, $T$, is much less trivial. One might expect that interference is suppressed when $T$ becomes larger than the level spacing, $\Delta.$ As was shown theoretically this is not the case neither for CI Jagla and Balseiro (1993); Dmitriev et al. (2010); Shmakov et al. (2012, 2013); Dmitriev et al. (2015, 2017) nor for HI Niyazov et al. (2018, 2020, 2021a, 2021b), and the interference effects “survive” even in the case
@@
where $L=L_{1}+L_{2},$ with $L_{1,2}$ the lengths of the interferometer arms, and $v_{\rm F}$ is the Fermi velocity. However, instead of resonances, $G(\phi)$ exhibits narrow antiresonances under the condition (1). Physically, antiresonances arise due to trajectories that interfere destructively at any energy $\varepsilon$ and, accordingly, are insensitive to energy averaging (see discussions in Dmitriev et al. (2010) and Niyazov et al. (2018) regarding CI and HI, respectively). For typical $v_{\rm F}$ of order of $10^{7}~{}$cm/s and typical system sizes ($>1$ micron), the value of $\Delta$ does not exceed several degrees Kelvin. This means that the interference effects can be studied at relatively high temperatures, which are relevant for various applications.
@@
The conductance in CI in the temperature range (1) has been studied in detail, in particular, considering the effects of the electron-electron interaction Dmitriev et al. (2010), disorder Shmakov et al. (2013), and spin-orbit interaction Shmakov et al. (2012). Recently, some of these results were generalized to HI systems: to single ring-shaped HI Niyazov et al. (2018, 2020, 2021a, 2021b) with a static magnetic defect and, partially, to arrays of coupled helical rings Niyazov et al. (2023).
@@
The influence of the interference effects on the noise has previously been discussed only in the low-temperature regime, $T\ll\Delta$, see review in Ref. Kobayashi and Hashisaka (2021) for CI and Refs.  Edge et al. (2013); Dolcini (2015) for HI. To the best of our knowledge, the effects of the interference for the noise in the high-temperature regime (1) were not discussed in detail for either CI or HI setups. On the other hand there is a clear need of the corresponding theory, which would allow direct comparison with experiment, since nanoscale rings made from conventional materials with a few or single conducting channels have been already fabricated [21–25]. Also, the experiment aimed to study helical interferometers was recently reported Munyan et al. (2023), although it focused on the low-temperature case. Note that in the latter experiment a quantum point contacts to the HES have been realized, thus allowing experimental verification of interference effects in conductance and noise of tunneling HI, especially having in mind that the condition (1) does not require very strict restrictions on temperature.
It is worth noting that the shot noise in HES has already been discussed for the case of infinite edge Lezmy et al. (2012); Del Maestro et al. (2013); Tikhonov et al. (2015); Nagaev et al. (2018); Väyrynen and Glazman (2017); Piatrusha et al. (2018); Kurilovich et al. (2019a, b); Pashinsky et al. (2020); Hsu et al. (2021); Probst et al. (2022), i.e. without taking into account the interference effects. In particular, a great attention was paid to infinite HES with a dynamic magnetic impurity Väyrynen and Glazman (2017); Kurilovich et al. (2019a, b); Pashinsky et al. (2020); Hsu et al. (2021); Probst et al. (2022)), whose magnetic moment changes direction after each scattering event. Discussing HI in this paper, we focus on the opposite case of static magnetic defect (see discussion of our model in the next section). The experimental measurement of the Fano factor of HI at zero magnetic field for the edge states of 2D TI gives the value $0.30.1<\mathcal{F}<$ 0.3  Tikhonov et al. (2015); Piatrusha et al. (2018). The upper value, 0.3, is close to the value $1/3$ for diffusive conductor. A similar result was obtained in the model with a large number of “islands”, allowing spin relaxation and tunnel-connected to the HES Aseev and Nagaev (2016). The effect of islands of different types on HES is currently being actively debated (see Olshanetsky et al. (2023) and references therein). However, the latter model, among others, does not consider interference effects.
@@
In this paper, we will discuss ballistic CI without impurities and defects and also HI with a single static magnetic defect (MD) placed into one of the arm of the interferometer. The goal of the work is to find the dependence of Fano factor, $\mathcal{F}$, on tunneling coupling, on the magnetic flux, and, in the case of HI, on the strength of magnetic defect. The latter is characterized by the parameter $\theta$, with $\theta=0$ in the absence of defect, and $\theta=\pi/2$ for the strong, fully reflective defect. We will find general analytical expressions for $\mathcal{F}^{\rm CI}$, $\mathcal{F}^{\rm HI}$ and analyze simple limiting cases assuming that inequality (1) is satisfied.
@@ II Model
We calculate the current shot noise in the Aharonov-Bohm interferometers of two types: conventional AB interferometer (CI) and helical AB interferometer (HI), depicted in Fig. 1a and Fig. 1b, respectively. They are based on spinless single-channel quantum wire (CI) and helical edge states (HI). We model leads by conventional single-channel wires (spinless or spinfull for the conventional and helical case, respectively). Although this model of the leads is very simplified, it is commonly used in quantum interferometry, starting from the work Büttiker et al. (1984), since, as we demonstrate below, it allows to describe the transition from the metallic leads to the tunnel ones. In particular, this model qualitatively describes quantum point contacts to the interferometer.
@@ II.1 Conventional AB interferometer
$\hat{S}^{\text{CI}}=$ $\begin{pmatrix}t_{\text{r}}&t_{\text{out}}&t_{\text{out}}\\ t_{\text{in}}&t_{\text{b}}&t\\ t_{\text{in}}&t&t_{\text{b}}\\ \end{pmatrix}\,,$ (2)
$t=$ $\frac{1}{1+\gamma}\,,\quad t_{\text{b}}=-\frac{\gamma}{1+\gamma}\,,$
$t_{\text{in}}=$ $t_{\text{out}}=\frac{\sqrt{2\gamma}}{1+\gamma}\,,\quad t_{\text{r}}=-\frac{1-\gamma}{1+\gamma}\,.$
More general case of $\hat{S}^{\text{CI}}$ is discussed in Refs. Aristov and Wölfle (2011) and Aristov et al. (2010). The scattering amplitudes entering Eq. (2) obey $12t_{\rm in}^{2}+t_{\rm r}^{2}=t^{2}+t_{\rm b}^{2}+t_{\rm out}^{2}=$ 1 . The meaning of these amplitudes is illustrated in Fig. 2: $t_{\rm in}$ is the amplitude to enter in the interferometer into one of the counter-propagating channels, $t_{\rm out}$ is the amplitude to exit the interferometer, $t$ and $t_{\rm b}$ are respectively the amplitudes of forward and backward scattering on the contact inside the interferometer, and $t_{\rm r}$ is the amplitude of backward scattering for electrons coming from the lead. All scattering amplitudes are parameterized by a single parameter $\gamma\in(0,\infty)$.
Figure 2: Scattering amplitudes entering scattering matrix Eq. (2) for CI. The lead is modeled by spinless 1D wire. For tunneling contact, $|t_{\rm r}|\approx 1,$ while the case of metallic contact corresponds to $t_{\rm r}=0.$
The meaning of the parameter $\gamma$ is further illustrated in Fig.   3, see also discussion in Refs. Dmitriev et al. (2010); Aristov et al. (2010, 2017). The case $\gamma\ll 1$ corresponds to tunneling contact, where almost closed interferometer is weakly coupled to the leads, with $\gamma$ being a tunneling transparency (see Fig. 3a). Interestingly, the case $\gamma\to\infty$ also corresponds to weak tunneling coupling. However, this case is strongly different from the case $\gamma\to 0,$ as illustrated in Fig. 3b. Metallic contact is modeled by $\gamma\approx 1$ (see Fig. 3c). It is worth noting that $t_{b}\neq 0$ for any non-zero $\gamma.$ Below we show that two configurations, Fig. 3a and Fig. 3b, may correspond to the same conductance but to different noise.
Figure 3: Schematic illustration of different types of leads: panels (a) and (b) shows two different types of tunneling coupling, corresponding to $\gamma=0$ and $\gamma=\infty,$ respectively; panel (c) illustrates “metallic” contact, $\gamma\approx 1.$

### v2

@@ Abstract
We study tunneling transport through quantum Aharonov-Bohm (AB) interferometers and demonstrate that interference effects strongly modify shot noise of the current. We discuss in detail two simplest setups: conventional single-channel spinless interferometer and interferometer formed by helical edge states of two-dimensional topological insulator. We demonstrate that both in the conventional and the helical case the interference dramatically changes the Fano factor and its magnetic field dependence. For weak tunneling coupling, the Fano factor of both setups exhibits a periodic series of sharp AB peaks depending on the magnetic flux piercing the system. Our key finding is that the Fano factor in the helical interferometer provides information about the presence of backscattering defects violating topological protection. In particular, the amplitude of AB peaks in the helical setup is proportional to the strength of the defect in contrast to conventional setup, where peaks have finite amplitude even in the ballistic case.
@@ I Introduction
Quantum interferometry is a rapidly growing area of electronics with a huge potential for practical use, in particular, in the booming fields of quantum computing and quantum cryptography. One of the beautiful manifestation of quantum interference is the Aharonov-Bohm effect Aharonov and Bohm (1959, 1963)—the sensitivity of the phase of electron wave function to magnetic flux. Quantum interferometers based on this effect are used for high-precision measurements of magnetic fields, for example, by using SQUID circuits.
@@
The simplest setup to study interference and particularly the Aharonov-Bohm effect is a ring-shaped interferometer with two single-channel arms placed in the perpendicular magnetic field. Due to the interference of electron trajectories winding around such a setup, the observables, like transmission coefficient and the noise intensity, show Aharonov-Bohm oscillations with a certain periods $\Delta\phi$ as a function of dimensionless magnetic flux $\phi=\Phi/\Phi_{0}$ threading the device, where $\Phi_{0}=hc/e$ is the flux quantum and $\Phi=BS,$ where $B$ is homogeneous magnetic field and $S$ is the area of the region encompassed by one-dimensional channel. The key statement of our paper is that noise in such a simple system is still not well understood and depends on specific realizations of the single-channel arms.
@@
Here, we will investigate and compare in details two types of setups: (i) a spinless single-channel interferometer based on conventional materials, i.e. a single-channel spinless quantum wire, folded in a ring-like shape and having two tunnel or metallic contacts attached (see Fig. 1a); (ii) an interferometer based on helical edge states. Below we refer to these systems as a conventional interferometer (CI) and a helical interferometer (HI), respectively. The latter interferometer can be implemented in two-dimensional topological insulators  Bernevig and Hughes (2013), which are insulators in the bulk, but have helical one-dimensional (1D) edge states conducting the current without dissipation Bernevig and Hughes (2013); Hasan and Kane (2010); Qi and Zhang (2011). These edge states are formed by electrons with opposite spins traveling in opposite directions, so that backscattering by non-magnetic impurities is prohibited. If one attaches two contacts to the edge and apply the gate voltage to shift the Fermi level into the band gap, then the conductance of such a device is fully determined by the edge states (see Fig. 1b). Since the sample boundary can be bypassed in two directions, possibly with several windings, such a system represents an interferometer that encodes information about properties of the helical edge states (HES).
Both setups, CI and HI, have two electron waves propagating clockwise and counterclockwise in 1D channel. However, the interference phenomena in these systems manifest themselves in a rather different way.
A spinless CI based on conventional materials (Fig. 1a) shows interference effects even in the simplest ballistic case, when both arms of interferometer do not contain any impurities. By contrast, the waves, propagating in opposite directions in ballistic HI, have also opposite spin projections and cannot interfere provided that contacts to the HES are non-magnetic (note that CI differs from Rashba-interferometers, where ballistic interference is possible due to the Aharonov-Casher effect Aronov and Lyanda-Geller (1993); Koga et al. (2004); Nitta and Bergsten (2007); Shmakov et al. (2012), see also Ref. Manchon et al. (2015) for review and references therein). Interference effects in HI appear in presence of at least one backscattering center (see Fig. 1b). It means that the study of interference effects can help to solve the well-known problem of the lack of topological protection in real samples (see recent experiment Olshanetsky et al. (2023) and discussion of the experimental situation therein). Specifically, observation of interference phenomena in the HI can give information about mechanisms of possible violation of the topological protection.
@@
Figure 1: Aharonov-Bohm interferometers based on conventional spinless single-channel quantum wire (a) and on helical edge channels of 2D topological insulator (b). The leads (shown by gray color) are modeled by conventional single-channel [spinless (a) and spinful (b)] wires. The black dot in panel (b) represents a backscattering impurity. Gray regions represent contacts described by $3\times 3$ and $4\times 4$ $S$-matrix for CI and HI, respectively [see Eqs. (2), (4) and Figs. 2, 4]. Homogeneous magnetic field, $B,$ is perpendicular to the picture plane. Corresponding magnetic flux, $\Phi=BS,$ is proportional to the area, $S,$ of the region encompassed by one-dimensional channels
@@
For both setups we will investigate the manifestations of quantum interference in the shot noise of the current. Such noise is a consequence of the discreteness of the electron charge. Importantly, its measurement provides information not available from conductance measurements: namely, the charge and statistics of current carriers, and the internal energy scales of the system de Jong and Beenakker (1997); Blanter and Büttiker (2000). Our primary goal is to calculate the Fano factor, $\mathcal{F}$, which is the ratio of the shot noise and the so-called Schottky noise value. This ratio is often used to characterize different transport regimes. For example, the values $\mathcal{F}=1/3$~ and $\mathcal{F}=1/2$~ for a diffusive conductor and tunneling transmission through a quantum level, respectively, are universal and independent of the system details de Jong and Beenakker (1997); Blanter and Büttiker (2000). As we demonstrate below, the Fano factor also is very sensitive to the interference effects which manifest itself in different way in CI and HI.
It is worth noting that the shot noise in HES has already been discussed for the case of infinite edge Lezmy et al. (2012); Del Maestro et al. (2013); Tikhonov et al. (2015); Mani and Benjamin (2017); Väyrynen and Glazman (2017); Piatrusha et al. (2018); Kurilovich et al. (2019a, b); Pashinsky et al. (2020); Hsu et al. (2021); Probst et al. (2022); Mishra and Benjamin (2023), i.e. without taking into account the interference effects. In particular, a great attention was paid to infinite HES with a dynamic magnetic impurity Väyrynen and Glazman (2017); Kurilovich et al. (2019a, b); Pashinsky et al. (2020); Hsu et al. (2021); Probst et al. (2022)), whose magnetic moment changes direction after each scattering event. Discussing HI in this paper, we focus on the opposite case of static backscattering defect, which is not necessarily magnetic (in particular, backscattering mechanisms involving charged puddles in the bulk of topological insulator are actively discussed Väyrynen et al. (2013)). We will describe such defect by $S$ matrix of general type. The experimental measurement of the Fano factor of HI at zero magnetic field for the edge states of 2D TI gives the value $0.1<\mathcal{F}<0.3$  Tikhonov et al. (2015); Piatrusha et al. (2018). The upper value, 0.3, is close to the value $1/3$ for diffusive conductor. A similar result was obtained in the model with a large number of “islands”, allowing spin relaxation and tunnel-connected to the HES Aseev and Nagaev (2016). The effect of islands of different types on HES is currently being actively debated (see also recent works Olshanetsky et al. (2023); Krainov et al. (2025) and references therein).
@@
Since the conductance and shot noise of interferometers of various types were also studied in a great number of works, we will start with a short description of the state of the art. Let us first recall the key results known for the conductance of the Aharonov-Bohm interferometer. At weak tunnel coupling and at low temperatures, the conductance $G(\phi)$ exhibits sharp resonances at certain values of magnetic flux, arising when quantum levels of the system are crossed by the Fermi level Büttiker et al. (1984, 1985).
As was understood later Jagla and Balseiro (1993), the case of relatively high temperatures, $T$, is much less trivial. One might suggest that interference is suppressed when $T$ becomes larger than the level spacing, $\Delta.$ However, as was first predicted in Ref. Jagla and Balseiro (1993) for CI, the conductance shows interesting interference-induced behavior in this case. Specifically, as it was shown theoretically, for both CI Jagla and Balseiro (1993); Dmitriev et al. (2010); Shmakov et al. (2012, 2013); Dmitriev et al. (2015, 2017) and HI Niyazov et al. (2018, 2020, 2021a, 2021b), the interference effects lead to resonance behavior of the conductance $G$ even in the case
@@
where $L=L_{1}+L_{2},$ with $L_{1,2}$ the lengths of the interferometer arms, and $v_{\rm F}$ is the Fermi velocity. However, instead of resonances, $G(\phi)$ exhibits narrow antiresonances under the condition (1). Physically, anti-resonances arise due to trajectories that interfere destructively at any energy $\varepsilon$ and, accordingly, are insensitive to energy averaging (see discussions in Dmitriev et al. (2010) and Niyazov et al. (2018) regarding CI and HI, respectively).
@@
The conductance in CI in the temperature range (1) has been studied in detail, in particular, considering the effects of the electron-electron interaction Dmitriev et al. (2010), disorder Shmakov et al. (2013), and spin-orbit interaction Shmakov et al. (2012). Recently, some of these results were generalized to HI systems: to single ring-shaped HI Niyazov et al. (2018, 2020, 2021a, 2021b) with a static magnetic defect and, partially, to arrays of coupled helical rings Niyazov et al. (2023). (See also discussion for other types of interferometers Feldman et al. (2007); Wang and Feldman (2010); Yang (2015).)
Surprisingly, non-trivial interference-induced characteristics of noise emerging under condition (1) were not discussed previously even for the simplest possible model of a single-channel CI, in view of the huge number of publications devoted to CI. To the best of our knowledge, the influence of the interference effects on the noise has previously been discussed only in the low-temperature regime, $T\ll\Delta$, see review in Ref. Kobayashi and Hashisaka (2021) for CI and Refs.  Edge et al. (2013); Dolcini (2015) for HI.
@@
On the other hand there is a clear need of the corresponding theory both for CI and HI, which would allow direct comparison with experiment. Indeed, nanoscale rings made from conventional materials with a few or single conducting channels have been already fabricated Shea et al. (2000); Piazza et al. (2000); Fuhrer et al. (2001); Keyser et al. (2003); Zou et al. (2007). Also, the experiment aimed to study HI was recently reported Munyan et al. (2023), although it focused on the low-temperature case. Note that in the latter experiment a quantum point contacts to the HES have been realized, thus allowing experimental verification of interference effects in conductance and noise of tunneling HI, especially having in mind that the condition (1) does not require very strict restrictions on temperature. Indeed, for typical $v_{\rm F}$ of order of $10^{7}$~cm/s and typical system sizes (>1>1 micron), the value of $\Delta$ does not exceed several degrees Kelvin. This means that the interference effects can be studied at relatively high temperatures, which are relevant for various applications. A comment about magnetic-field-induced phenomena is in order to do. Actually, the magnetic field causes two effects: (i) the AB phase appears and (ii) the topological protection is destroyed. Simple estimates show that the flux quantum, $\Phi\sim\Phi_{0}=hc/e,$ is achieved for sample with micron size in fields $B\sim 3$ Oe, well below the expected magnitude of the fields destroying the edge states Du et al. (2015); Zhang et al. (2014); Hu et al. (2016).
Recently we discussed shot noise in HI Niyazov et al. (2024). We considered HI with identical contacts and calculated the noise power in the regime (1) paying attention to the interference effects. We found the Fano factor strongly depending on the tunneling coupling and strength of backscattering defect.
Here, we further develop the results obtained in Ref. Niyazov et al. (2024). We present a detailed study of the sharp Aharonov-Bohm peaks in the Fano factor of the noise. We also demonstrate that the simultaneous measurement of the conductance and the Fano factor of HI allows one to directly determine the strength of the backscattering defect even in the absence of detailed information about scattering amplitudes characterizing the contacts. Having also in mind the recent experiment Munyan et al. (2023), we generalize here the calculation of the Fano factor of HI for experimentally relevant case of non-equivalent contacts.
In Ref. Niyazov et al. (2024), we also briefly discussed the case of CI, announcing without derivation the formula for the Fano factor of the CI with symmetric arms and identical contacts. In this paper we provide detailed calculations, and generalize our analysis for the experimentally relevant case of asymmetric arms and non-identical contacts Shea et al. (2000); Piazza et al. (2000); Fuhrer et al. (2001); Keyser et al. (2003); Zou et al. (2007). We study in detail the dependence of the noise of a CI on the magnetic field. It turns out that the interference effects are much more pronounced for CI and, consequently, the influence of the magnetic field on the Fano factor is much stronger as compared to HI. The reason is that the interference comes into play for CI already in the ballistic case, when impurities in both arms of CI are absent, while the Aharonov-Bohm peaks in HI are proportional to the backscattering probability.
@@
We will discuss ballistic CI without impurities and defects and also HI with a single static defect (which can be magnetic or non-magnetic) placed into one of the arm of the interferometer. The goal of the work is to find the dependence of Fano factor, $\mathcal{F}$, on tunneling coupling, on the magnetic flux, and, in the case of HI, on the strength of backscattering defect. We will find general analytical expressions for $\mathcal{F}^{\rm CI}$, $\mathcal{F}^{\rm HI}$ and analyze simple limiting cases assuming that inequality (1) is satisfied.
@@ II Model
We calculate the current shot noise in the Aharonov-Bohm (AB) interferometers of two types, CI and HI, depicted in Fig. 1a and Fig. 1b, respectively. They are based on spinless single-channel quantum wire and helical edge states. We model leads by conventional single-channel wires (spinless or spinful for the conventional and helical case, respectively). Although this model of the leads is very simplified, it is commonly used in quantum interferometry, starting from the work Büttiker et al. (1984), since, as we demonstrate below, it allows describing the transition from the metallic leads to the tunnel ones. In particular, this model qualitatively describes quantum point contacts to the interferometer.
@@ II.1 Conventional AB interferometer
$\hat{S}^{\mathrm{CI}}=$ $\begin{pmatrix}t_{\mathrm{r}}&t_{\mathrm{out}}&t_{\mathrm{out}}\\ t_{\mathrm{in}}&t_{\mathrm{b}}&t\\ t_{\mathrm{in}}&t&t_{\mathrm{b}}\\ \end{pmatrix}\,$, (2)
$t=$ $\frac{1}{1+\gamma}\,,\quad t_{\mathrm{b}}=-\frac{\gamma}{1+\gamma}\,$,
$t_{\mathrm{in}}=$ $t_{\mathrm{out}}=\frac{\sqrt{2\gamma}}{1+\gamma}\,,\quad t_{\mathrm{r}}=-\frac{1-\gamma}{1+\gamma}\,.$
More general case of $\hat{S}^{\mathrm{CI}}$ is discussed in Refs. Aristov and Wölfle (2011) and Aristov et al. (2010). The scattering amplitudes entering Eq. (2) obey $2t_{\rm in}^{2}+t_{\rm r}^{2}=t^{2}+t_{\rm b}^{2}+t_{\rm out}^{2}=1.$ The meaning of these amplitudes is illustrated in Fig. 2: $t_{\rm in}$ is the amplitude to enter the interferometer into one of the counter-propagating channels, $t_{\rm out}$ is the amplitude to exit the interferometer, $t$ and $t_{\rm b}$ are respectively the amplitudes of forward and backward scattering on the contact inside the interferometer, and $t_{\rm r}$ is the amplitude of backward scattering for electrons coming from the lead. All scattering amplitudes are parameterized by a single parameter $\gamma\in(0,\infty)$.
Figure 2: Scattering amplitudes entering scattering matrix Eq. (2) for CI. The lead is modeled by spinless 1D wire. For tunneling contact, $|t_{\rm r}|\approx 1,$ while the case of metallic contact corresponds to $t_{\rm r}=0.$ Contact region with area $S_{0}$ is shown by gray color. We assume that $BS_{0}\ll\Phi_{0},$ so that the magnetic field does not affect S-matrix.
The meaning of the parameter $\gamma$ is further illustrated in Fig.  3, see also discussion in Refs. Dmitriev et al. (2010); Aristov et al. (2010, 2017). The case $\gamma\ll 1$ corresponds to tunneling contact, where almost closed interferometer is weakly coupled to the leads, with $\gamma$ being a tunneling transparency (see Fig. 3a). Interestingly, the case $\gamma\to\infty$ also corresponds to weak tunneling coupling. However, this case is strongly different from the case $\gamma\to 0,$ as illustrated in Fig. 3b. Metallic contact is modeled by $\gamma\approx 1$ (see Fig. 3c). The scattering matrices describing these three cases are given by (see Eq. (2)):
$\hat{S}^{\mathrm{CI}}(\gamma=0)=$ $\begin{pmatrix}-1&0&0\\ 0&0&1\\ 0&1&0\\ \end{pmatrix}\,,\,\hat{S}^{\mathrm{CI}}(\gamma=\infty)=$ $\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&-1\\ \end{pmatrix}\,$, (3)
$\hat{S}^{\mathrm{CI}}(\gamma=1)=$ $\begin{pmatrix}0&\sqrt{2}&\sqrt{2}\\ \sqrt{2}&-1&1\\ \sqrt{2}&1&-1\\ \end{pmatrix}/2\,.$
It is worth noting that $t_{b}\neq 0$ for any non-zero $\gamma.$ Below we show that two configurations, Fig. 3a and Fig. 3b, may correspond to the same conductance but to different noise.
Figure 3: Schematic illustration of different types of leads (contact area is shown by gray color): panels (a) and (b) shows two different types of tunneling coupling, corresponding to $\gamma=0$ and $\gamma=\infty,$ respectively; panel (c) illustrates “metallic” contact, $\gamma\approx 1.$ Arrows in panels (a) and (b) show processes with non-zero amplitudes (compare with Fig. 2). Panel (c) corresponds to a “metallic contact”, so that all processes in Fig. 2 are allowed. Scattering matrices, describing cases (a), (b), and (c) are given by Eq. (3).